OpenStudy (anonymous):

What is the property of real numbers illustrated by this equation? A. Commutative Property of Multiplication B. Associative Property of Multiplication C. Associative Property of Addition D. Distributive Property

OpenStudy (anonymous):

\[(2+\sqrt{5})+3=2+(\sqrt{5}+3)\]

OpenStudy (anonymous):

I think it might be C

OpenStudy (anonymous):

thts wat i thought too, thanks :D

OpenStudy (anonymous):

No problem good luck

OpenStudy (anonymous):

well its either that or distributive property, im not sure

OpenStudy (anonymous):

Im pretty sure its C:) lol, if not im sorry

OpenStudy (anonymous):

ok thanks for trying

OpenStudy (anonymous):

No probs:)